Assume you are given the following relationships for the Megalops Corporation:

Sales/Total assets 1.6

Returns on Assets 7%

Return on Equity 15%

A. Calculate Megalops profit margin

B. Calculate Megalops Liabilities-To-Assets ratio.

C) Suppose 60% of its liabilities are in the form of debt. Calculate the debt-to-assets ratio.
Answer #1

A. Megalops Profit Margin =Return on Assets /(Sales/Total Assets) =7%/(1.6) =4.375% or 4.38%
B Megalops Liabilities-To-Assets ratio =1-(Return on Assets)/Return on Equity =1-7%/15% =8/15 =0.53
C. Debt to Assets ratio =60%*8/15 =0.32